# Hadwiger Models: Low-Temperature Behavior in a Natural Extension of the Ising Model

### v1

@@ Abstract
All isometrically invariant Markov (strictly local) fields on binary assignments are induced by energy functions that can be represented as linear combinations of area, perimeter, and Euler characteristic. This class of model includes the Ising model, both ferro- and antiferro-magnetic, with and without a field, as well as the “triplet” Ising model We determine the low-temperature behavior for this class of model, and construct a phase diagram of that behavior. In particular, we identify regions with three geometric phases, regions with a single unique phase, and coexistence lines between them.
@@ 1 Introduction
Figure 1: Converting binary assignments to polyconvex subsets
@@
Traditionally, the Ising Hamiltonian is formally defined as a map $H:\{-1,1\}^{V}\to\mathbb{R}$ by $H(\sigma)=\sum_{(x,y)\in E}\sigma(x)\sigma(y)+h\sum_{x\in V}\sigma(x)$, where $V,E$ are the vertex and edge sets of some graph $G$. However, the phase transition of the two-dimensional Ising model was first proved by Peierls [P36] using the equivalent formulation where $H:\{0,1\}^{F^{*}}\to\mathbb{R}$ by $H=P(\sigma)+hA(\sigma)$ where $F^{*}$ are the faces of the dual graph $G^{*}$ of $G$, and $P,A$ are respectively the perimeter and area of $\cup_{f\in\sigma}f$, as represented in figure 1. This formulation allowed Peierls to use a geometric argument to bound the perimeter and thus guarantee that one orientation dominates when the temperature is sufficiently low.
@@
Perimeter and area are both invariant valuations, that is, given a union of convex sets (polyconvex) $U$, they satisfy $f(T(U))=f(U)$ for any isometry $T$, $f(U\cup V)=f(U)+f(V)-f(U\cap V)$, and $f(\varnothing)=0$. $f(U\cup V)=f(U)+f(V)-f(U\cap V)$ if and only if the energy depends only on immediately neighboring sites, as repeated application of the equation decomposes the energy function into a sum over faces, edges, and vertices. This decomposition property means the distribution within a given region depends only on its boundary; such a distribution is called a Markov field. For an infinite graph with finite fundamental domain, the dual faces are always compact and polygonal, so their union is polyconvex, and any isometrically invariant Markov Hamiltonian on binary assignments is a valuation, up to a constant term for the energy of the empty set. Hadwiger’s theorem describes the set of all valuations explicitly:
Theorem (Hadwiger’s Theorem[H56]).
@@
Any invariant valuation must be a linear combination of the perimeter, area, and Euler characteristic
@@
Hadwiger’s theorem implies that any invariant Markov Hamiltonian must be a linear combination of perimeter term, area term, and Euler characteristic. In fact, all Markov fields are induced by Markov Hamiltonians [G73], so this class of function includes every possible Markov field on binary assignments. This paper is dedicated to the study of the distributions induced by members of this class of Hamiltonian on the hexagonal lattice. Note that the Euler characteristic can be expressed as a sum over particular subsets $S$ of $\prod_{f\in S}\sigma_{f}$, so this class is a special case of the generalized Ising models. We choose to work on the hexagonal lattice because each term behaves nicely under inversion: if a set of faces is replaced with its complement, the perimeter remains unchanged, whereas the area and the Euler characteristic are inverted, up to an additive constant.
@@
Define a domain as a connected subset $F$ of the faces of some lattice with connected complement, a configuration $\sigma\in\{0,1\}^{F}$ on a domain as a binary assignment to those faces, a boundary configuration $\rho$ as a binary assignment to all the faces in the lattice, an occupied or filled hexagon $f$ as one such that $\sigma(f)=1$ if $f\in\Lambda$ or $\rho(f)=1$ if $f\notin\Lambda$ and, $\partial\Lambda$ (the “boundary”) as the set of faces neighboring $\Lambda$. We identify $\sigma$ and $\rho$ interchangeably with the associated set of occupied hexagons.
Definition (Hadwiger Model).
@@
Given a domain $\Lambda$, and a boundary configuration $\rho$, at a temperature $T$, define a probability distribution on the set of configurations $\Omega$ with parameters $x,p,a$ by
$\mu(\sigma)\coloneqq\mu_{x,p,a,T}(\sigma)\propto e^{\mathcal{H}(\sigma)/T},\;\mathcal{H}=x\chi(\sigma)+pP(\sigma)+aA(\sigma)$
@@
While we define the Hadwiger models in terms of the Ising, area, and Euler functions, the phase diagram is naturally parametrized by an equivalent formulation in terms of vertex energies. Consider the energy as a sum over vertices, and evaluate each term in the Hamiltonian on each of the four possible vertex states. The associated energy assignments are represented in figure 3. Note that only number and not position of adjacent hexagons matters, because each model is rotation invariant, and that every model assigns the empty state energy 0. Refer to the vertex state with 0,1,2, and 3 incident filled hexagons as $E,C,H,F$ respectively. We can notate the energy assigned by the Hamiltonian to vertex state $X$ with $e_{X}$.
@@
Proposition.
Area, perimeter, and Euler characteristic span the full space of possible energy assignments to vertex states with $e_{E}=0$, so the set of Hadwiger functions is
$\{g(\sigma)=\sum_{v\in V}f(\sigma_{v})|f(\sigma_{v})=e_{C}\cdot\delta_{C,\sigma_{v}}+e_{H}\cdot\delta_{H,\sigma_{v}}+e_{F}\cdot\partial_{F,\sigma_{v}}\}$
@@
where $\sigma_{v}$ is the vertex state induced by $\sigma$ on $v$, and $\delta$ is the Kronecker delta.
@@
Coordinates of $e_{C},e_{H},e_{F}$ are the most natural parametrization to consider when evaluating low temperature behavior, as they determine the ground configuration. In particular, the ground configurations are those which consist only of the lowest-energy vertex states for that model. The full space of Hamiltonians is 3-dimensional, but two Hamiltonians related by a scaling factor are considered to represent the same model at different temperatures. Therefore, the space of all “models” is a sphere. If we unwrap this sphere to the plane, we have a 2 dimensional diagram of the full set of models, represented in Figure 4.
@@
Figure 4: At $T=0$, the behavior is determined only by states with minimal energy. Along three of the transition lines, where multiple vertex states have minimal energy, there is nonzero entropy, and so these represent “models” in a fuller sense than other points, where only finitely many configurations are permitted. When both $C$ and $H$ vertex states are minimal, the allowed configurations are those of the 0 temperature antiferromagnetic Ising model, which are the same as those of the dimer model [BH82]. When both $E$ and $C$ vertex states are minimal, the allowed configurations are those of the hard hexagon model, with the even distribution on allowed configurations [B89]. The $F$ and $H$ line is symmetric to the $E$ and $C$ under inversion. “Landmark points” are marked, including the pure Euler, Perimeter, and Area terms.
@@
Each region in the diagram is characterized by the vertex state with the lowest energy, with the transition lines representing points where two vertex states have equal and minimal energy. Continuing along transition lines maintains the equality of the energies even when they are no longer minimal, so their intersections represent particularly symmetrical points. The $H$ region has three ground configurations, each having every vertex incident to two hexagons, and possessing maximally many holes. Each ground configuration corresponds to one of the 3 sublattices in the hexagonal lattice. The $C$ region is the inverse of the $H$ region, having three ground configurations with maximally many components. The $F$ and $E$ regions have one ground configuration each, such that the value on every hexagon is 1 and 0 respectively. You can see the structure of each in Figure 3
2 Main Result
@@
First, we will recall some terms, so we can use them freely in the following sections. A Gibbs state of a model with Hamiltonian $\mathcal{H}$ is a mapping from bounded measurable functions $f$ on the set of configurations $\Omega$ to expectations $\langle f\rangle$ of those functions, such that for any finite region, the expectations conditional on a particular boundary configuration $\rho$ are those generated in the above way by the Hamiltonian, given $\rho$. Note that this implicitly defines probability distributions on configurations for any finite regions via indicator functions. Such a Gibbs state is “extremal” if it cannot be represented as a linear combination of other Gibbs states with positive coefficients. A configuration or distribution is “translation-invariant” if it is invariant under all translations that preserve the lattice, while it is “periodic” if it’s invariant spanning sublattice. These terms are technically distinct, but any periodic model can be rendered translation-invariant by considering larger configuration spaces on tiles formed from the fundamental domain of the period lattice [D85]. The Peierls condition states that there exists a finite collection of translation-invariant minimal-energy configurations and a constant $c$ such that, for any other configuration $\omega$ that differs from a ground configuration $\eta$ only on a finite set $\Gamma$, $\mathcal{H}(\omega)-\mathcal{H}(\eta)\geq c|\Gamma|$. A transition line refers to a curve such that multiple configurations, not related by a symmetry, have minimal energy. Conversely, a coexistence line refers to a curve such that there is more than one Gibbs state, which are not related by a symmetry.
@@
With the space of models laid out, we now consider the low temperature behavior of the Hadwiger models. There is a chain of statements we must string together to restrict our Gibbs states. Dobrushin tells us that in 2-dimensions and at low temperatures, all Gibbs states of a translation-invariant (resp. periodic) Peierls model are translation-invariant (periodic)[D85]. Zahradnik, extending Pirogov–Sinai, tells us that, given the Peierls condition, all translation invariant (or equivalently, periodic) Gibbs states are linear combinations of extremal “pure phases”, each of which is dominated by a particular ground configuration of the model [Z84]. Slawny tells us the position of the transition curves between areas such that different ground configurations have an extremal Gibbs state corresponding to them for a particular model . [S87]
Definition.
@@
Given a ground configuration $G$, notate by $B(\omega)$ the set of vertices not agreeing with $G$ We refer to a Gibbs state as “dominated” by a certain global configuration $G$, or equivalently as a “pure phase” if the portion of vertices in $B(\omega)$ goes to 00 as $T\to 0$, and every connected component of $B(\omega)$ is finite. [PS75]. These pure phases are exactly those induced by choosing boundary conditions agreeing with that configuration, although they may be induced by other boundary conditions. To respect the symmetry of the model, if a set of $n$ configurations is symmetrically related under a given model, we instead mean that the model has a $1/n$ chance of being dominated by any of the $n$ pure Gibbs states given general boundary conditions.
@@
In the interior of any region such that a single configuration (or symmetric set of configurations) has minimal energy density, there will always be a temperature low enough that the model is dominated by the lowest energy configurations. However, at any fixed low temperature, the lowest energy configuration may not dominate in the entire region. The asymptotic transition curves between areas where different phases dominate, which also determine the behavior at any nonzero low temperature of models with more than one distinct asymmetric ground configuration, can be calculated based on the technique of Slawny [S87]. Using these techniques, we generate a phase diagram representing the low-temperature behavior of the Hadwiger models, depicted in Figure 5
@@
Figure 5: Phase diagram showing the regions where given configurations dominate. Each color, including white, represents a region where a particular set of configurations dominates. On the boundaries between the regions, multiple Gibbs states not related by a symmetry dominate. Note that the dark red region surrounding the non-Peierls transition lines indicates points where Pirogov–Sinai techniques are not applicable. This region shrinks with decreasing temperature, but is never empty. On some of the non-Peierls lines, we use disagreement percolation to prove no domination at any temperature.
3 Proving the Phase Diagram

### v2

@@ Abstract
In two dimensions, all isometrically invariant Markov random fields on binary assignments are induced by energy functions that can be represented as linear combinations of area, perimeter, and Euler characteristic. This class of model includes the Ising model, both ferro- and antiferromagnetic, with and without a field, as well as the Baxter-Wu model. On the hexagonal lattice, we determine the low-temperature behavior for this class of model, and construct a phase diagram of said behavior. In particular, we identify regions with three geometric phases, regions with a single unique phase, and coexistence curves between them. We also characterize the behavior along two non-Peierls lines, where entropy fails to vanish the as temperature goes to zero.
keywords:
Lattice statistical mechanics, Spin models, Hadwiger’s theorem, Ising model
@@ 1 Introduction
Traditionally, the Ising Hamiltonian is defined as a function $H:\{-1,1\}^{V}\to\mathbb{R}$ by $H(\sigma)=\sum_{(x,y)\in E}\sigma(x)\sigma(y)+h\sum_{x\in V}\sigma(x)$, where $V$ and $E$ are the vertex and edge sets of some graph $G$. However, existence of a phase transition for the planar Ising model was first proven by Peierls [P36] using an equivalent formulation where spins are assigned to the faces $F$ of the dual graph $G^{*}$ of $G,$ and $H$ is defined in terms of the area and perimeter of the union of dual faces assigned spin $1$. Formally, given a spin assignment $\sigma\in\{-1,1\}^{F}$ denote by $\sigma^{*}$, represented in Fig. 1, the collection of dual faces assigned spin $1$ and set $H(\sigma^{*})=P(\sigma^{*})+hA(\sigma^{*})$, where $P(\sigma^{*})$ and $A(\sigma^{*})$ are respectively the area and perimeter of $\sigma.$ This formulation allowed Peierls to bound the perimeter weights with a geometric argument and thus, at $h=0$, guarantee that one orientation dominates at sufficiently low temperatures.
Figure 1: Converting binary assignments on a lattice to polyconvex subsets on the faces of the dual lattice. $\sigma^{*}$ is illustrated by the union of the black hexagons on the right.
@@
A valuation is a function $H$ assigning real numbers to polyconvex subsets of $\mathbb{R}^{d}$ (where a polyconvex set is one that is a finite union of bounded convex sets) such that $H(\varnothing)=0$ and $H(U\cup V)=H(U)+H(V)-H(U\cap V)$. Functions satisfying the latter equation have the Markov property. If $H$ also satisfies $H(T(U))=H(U)$ for all isometries $T$, then $H$ is called an invariant valuation. The Markov property implies that the Gibbs distribution induced by $H$ within a given region depends only on its boundary; if the configuration space assigns positive probability to all possible assignments from the set to the lattice, such a distribution is called a Markov random field. For an infinite planar graph with finite fundamental domain, the dual faces are always compact and polygonal, meaning their union is polyconvex, so any isometrically invariant Markov Hamiltonian on binary assignments is an invariant valuation, up to a constant term for the energy of the empty set. Hadwiger’s theorem describes the set of all such functions explicitly:
Theorem 1 (Hadwiger’s Theorem[H56]).
@@
In two dimensions, every invariant valuation must be a linear combination of perimeter, area, and Euler characteristic.
Recall that in two dimensions Euler characteristic $\chi$ can be expressed as
$\chi=\#\text{vertices}-\#\text{edges}+\#\text{faces}.$
All Markov random fields are induced by Markov Hamiltonians [CH71, G73], so the class of Gibbs distributions induced by invariant valuations coincides with the class of Markov random fields on binary assignments.  [^1] The relation between Markov Hamiltonians and Markov random fields fails to hold without the positivity condition: for extensions of the statement, see [CM16]. We summarize the preceding discussion in a corollary.
Corollary 2.
A Markov random field in two dimensions is the Gibbs distribution induced by a Hamiltonian that is a linear combination of perimeter, area, and Euler characteristic.
Note that each term can be expressed in terms of a product over a set of spins, so this class forms a subfamily within the generalized Ising models. Hadwiger’s theorem and the previous discussion extend to $d$-dimensions, where perimeter, area, and Euler characteristic are replaced by a collection of $d+1$ distinguished functions called intrinsic volumes. See [G07] for a definition. We leave the study of Markov random fields in $d$-dimensions to future work.
@@
This paper is dedicated to the study of the distributions induced by members of this class of Hamiltonian on the hexagonal lattice. We choose to work on the hexagonal lattice because each term behaves nicely under a spin flip: if a set of faces is replaced with its complement, the perimeter remains unchanged, whereas the area and the Euler characteristic are inverted, up to an additive constant.
@@
We introduce some notation. Define a domain as a subset $\Lambda$ of the faces of a lattice such that $\cup_{f\in\Lambda}f$ and $\mathbb{R}^{2}\setminus\cup_{f\in\Lambda}f$ are both connected, a configuration $\sigma\in\{-1,1\}^{\Lambda}$ on a domain as a binary assignment to the faces of $\Lambda$, a boundary configuration $\rho$ as a binary assignment to all the faces in the lattice, an occupied or filled hexagon $f$ as one such that $\sigma(f)=1$ if $f\in\Lambda$ or $\rho(f)=1$ if $f\notin\Lambda$ and, $\partial\Lambda$ (the “boundary”) as the set of faces neighboring $\Lambda$. We identify $\sigma$ and $\rho$ with their associated set of occupied hexagons.
Definition 1 (Hadwiger Model).
@@
Given a domain $\Lambda$, a boundary configuration $\rho$, and a temperature $T$, define a probability distribution on the set of configurations $\Omega=\{-1,1\}^{\Lambda}$ with parameters $x,p,a$ by
$\mu(\sigma)\coloneqq\mu_{x,p,a,T}(\sigma)\propto e^{\mathcal{H}(\sigma)/T},\;\mathcal{H}(\sigma)=x\chi(\sigma)+pP(\sigma)+aA(\sigma)$ (1)
@@
While we define the Hadwiger models in terms of the perimeter, area, and Euler functions, the phase diagram is naturally parametrized by vertex energies. As the Hamiltonian is a sum of interactions between neighboring hexagons, we can rewrite it as a sum over the vertices of functions depending only on the states of neighboring hexagons. Note that every model assigns the empty state energy 0, and that only number and not position of adjacent hexagons affect the energy, because the Hamiltonian is invariant under rotations. Refer to the vertex states with 0,1,2, and 3 incident filled hexagons as $E,C,H,F$ respectively. We denote the energy assigned by the Hamiltonian to vertex state $X$ with $e_{X}$. Assuming translation and rotation invariance, these energies are uniquely determined by the equation $\mathcal{H}=\sum_{v\in V}e_{X\left(v\right)}.$
Euler characteristic, perimeter, and area are signed sums over the number of adjacent faces, edges, or vertices, so for each element we sum over, we can divide by the number of vertices incident to that element to determine the energy of each vertex in a particular state. For example, the $H$ state vertex has two incident hexagons, three incident edges, and one incident vertex, so it contributes $1=\frac{2}{2}$ to the perimeter, $\frac{1}{3}=\frac{2}{6}$ to the area, and $\frac{2}{6}-\frac{3}{2}+1=-\frac{1}{6}$ to the Euler characteristic. Applying this logic to every state, we determine the vertex energies are
$e_{C}=\frac{x}{6}+p+\frac{a}{6},\quad e_{H}=-\frac{x}{6}+p+\frac{a}{3},\quad e_{F}=\frac{a}{2}\,.$ (2)
These assignments, represented in Fig. 3, span all choices of $e_{C},e_{H},e_{F}$, so any choice will be an invariant valuation. Combining this with the previous corollary yields the next result.
@@
Proposition 1.
The set of Markov Hamiltonians on the hexagonal lattice is
$\{g(\sigma)=\sum_{v\in V}f(\sigma_{v})|f(\sigma_{v})=e_{C}\cdot\delta_{C,\sigma_{v}}+e_{H}\cdot\delta_{H,\sigma_{v}}+e_{F}\cdot\delta_{F,\sigma_{v}}\}$ (3)
@@
where $\sigma_{v}$ is the vertex state induced by $\sigma$ on $v$ and $\delta$ is the Kronecker delta function.
In particular, we can obtain the Hamiltonian with parameters $x,p,a$ by using the vertex energies given in equation 2. On the other hand, we can convert back to $x,p,a$ parameters above with the equations
$x=3e_{C}-3e_{H}+e_{F},\quad p=\frac{e_{C}+e_{H}-e_{F}}{2},\quad a=2e_{F}.$
Similar reparametrizations exist for Markov Hamiltonians on other lattices and in higher dimensions, but they won’t necessarily include all possible energy assignments to vertex states.
@@
Because they determine the ground configuration, the coordinates $e_{C},e_{H},e_{F}$ are the most natural parametrization at low temperatures. In particular, the ground configurations, depicted in Fig. 3, are those which consist only of the lowest-energy vertex states for that model. The full space of Hamiltonians is 3-dimensional, but two Hamiltonians related by a scaling factor represent the same model at different temperatures, so the space of all “models” is a sphere. If we unwrap this sphere to the plane, we obtain a two dimensional diagram depicted in Fig. 4.
@@
Figure 4: At $T=0$ the behavior is determined only by states with minimal energy. Along three of the degeneracy lines where multiple vertex states have minimal energy there is nonzero entropy, meaning infinitely many global states are minimal. When both $C$ and $H$ vertex states are minimal the allowed configurations are those of the $T=0$ antiferromagnetic Ising model, the same as those of the dimer model [BH82]. When both $E$ and $C$ vertex states are minimal the allowed configurations are those of the hard hexagon model with the even distribution on allowed configurations [B89]. The $F$ and $H$ line is symmetric to the $E$ and $C$ under a spin flip. “Landmark points” are marked, including the pure Euler, Perimeter, and Area terms. The line containing “Pure Ferro Ising,” “Pure Area”, and “Pure Anti-Ferro Ising” corresponds to the Ising model with external field.
@@
The diagram decomposes into regions where there is a unique vertex state with lowest energy, separated by degeneracy lines where two vertex states have equal and minimal energy. The $H$ region has three ground configurations, corresponding to the three sublattices of the hexagonal lattice, which maximize the number of holes. The $C$ region is the inverse of the $H$ region, having three ground configurations with maximally many components. The $F$ and $E$ regions have one ground configuration each, including all and none of the hexagons respectively.
2 Main Results
@@
First, we will recall some terms so we can use them freely in the following sections. A Gibbs state of a model with Hamiltonian $\mathcal{H}$ is a mapping from bounded measurable functions $f$ on the set of configurations $\Omega$ to expectations $\langle f\rangle$ of those functions, such that for any finite region the expectations conditional on a particular boundary configuration $\rho$ are those generated by the Hamiltonian given $\rho$, as defined in Definition 1. Note that this implicitly defines probability distributions on configurations for any finite regions via indicator functions. Such a Gibbs state is “extremal” if it cannot be represented as a linear combination of other Gibbs states with positive coefficients. A configuration or distribution is “translation-invariant” if it is invariant under all translations that preserve the lattice, while it is “periodic” if it is invariant under a spanning sublattice. These terms are technically distinct, but any periodic model can be rendered translation-invariant by considering larger configuration spaces on tiles formed from the fundamental domain of the period lattice [DS85]. A model satisfies the Peierls condition if there exists a finite collection of translation-invariant minimal-energy configurations and a constant $c$ such that, for any other configuration $\omega$ that differs from a ground configuration $\eta$ only on a finite set $\Gamma$, $\mathcal{H}(\omega)-\mathcal{H}(\eta)\geq c|\Gamma|$.
Definition 2.
A degeneracy line refers to a curve such that multiple configurations, not related by a symmetry, have minimal energy, while a coexistence curve is a curve on which there are multiple non-symmetric Gibbs states.
@@
With the space of models laid out, we now consider the low temperature behavior of the Hadwiger models. There is a chain of statements we must string together to restrict our Gibbs states. Dobrushin and Shlosman tell us that in two dimensions and at low temperatures, all Gibbs states of a translation-invariant (resp. periodic) Peierls model are translation-invariant (periodic)[DS85]. Pirogov and Sinai tell us that given the Peierls condition, all translation invariant (or equivalently, periodic) Gibbs states are linear combinations of extremal “pure phases” [PS75]. Zahradnik, extending Pirogov–Sinai, tells us these “pure phases” are dominated by a particular ground configuration of the model [Z84]. Slawny and Bricmont tell us the position of the transition curves between areas such that different ground configurations have an extremal Gibbs state corresponding to them for a particular model [BS89][S87].
Definition 3.
@@
Given a ground configuration $G$ denote by $B(\omega)$ the set of vertices not agreeing with $G.$ We say that a Gibbs state is “dominated” by a certain global configuration $G$ — or equivalently is a “pure phase” — if there exists some constant $c$ such that the probability a vertex is in a cluster of vertices of perimeter $\Gamma$ not agreeing with $G$ is bounded above by $e^{-c\Gamma}$ [Z84]. These pure phases are exactly those induced by choosing boundary conditions agreeing with that configuration, although they may be induced by other boundary conditions. To respect the symmetry of the model, if a set of $n$ configurations is symmetrically related, we instead mean that the model has a $1/n$ chance of being dominated by any of the $n$ pure Gibbs states, conditional on an even measure for all boundary conditions.
@@
For any point in the interior of a region such that a single vertex state has minimal energy (that is, the pink, orange, blue and white regions in Fig. 4), there will be a temperature low enough that the minimal configuration dominates. However, at any fixed low temperature, the lowest energy configuration may not dominate in the entire region. The asymptotic transition curves between areas where different phases dominate, along which more than one distinct asymmetric ground configuration dominates, can be calculated based on the technique of Slawny [S87][BS89]. Using these techniques, we derive a phase diagram representing the low-temperature behavior of the Hadwiger models, depicted in Fig. 5 and summarized in the following theorems:
Theorem 3.
At any point such that the unique lowest energy vertex configuration is $H$ or $C$, there exists $T_{0}$ such that for $T<T_{0}$ there are exactly three distinct extremal Gibbs states, dominated respectively by the three ground configurations, and every Gibbs state is a linear combination of these.
To identify the position of the coexistence curves relative to the degeneracy lines, we determine which state dominates along the degeneracy lines, which in turn tells us where the line and curve cross.
Theorem 4.
For a fixed closed subset of the EHE-H degeneracy line not containing its endpoints there exists $T_{0}$ such that, $\forall 0<T<T_{0}$, if $e_{F}>e_{C}$ the $E$ state dominates and if $e_{F}<e_{C}$ the $H$ state dominates.
For example, along the closed subset $\min(E_{C},E_{F})\geq\epsilon$ of the degeneracy line $e_{H}=e_{E}=0$, there is a fixed $T_{0}$ such that the $E$ state dominates when $E_{C}<E_{F}$ and the $H$ state dominates when $E_{C}>E_{F}$. Moreover, the coexistence curve between the $E$ and $H$ states crosses this degeneracy line at the point $e_{F}=e_{C}=\frac{\sqrt{2}}{2},e_{H}=e_{E}=0$, passes inside the region where $H$ is the ground state when $e_{F}>e_{C},$ and is contained inside the region where $E$ is the ground state when $e_{F}<e_{C}.$
The corresponding statement under a spin flip also holds for the CFC-F degeneracy line.
Theorem 5.
For a fixed closed subset of the EFE-F degeneracy line not containing its endpoints there exists $T_{0}$ such that, $\forall 0<T<T_{0}$, if $e_{H}>e_{C}$ the $E$ state dominates and if $e_{H}<e_{C}$ the $F$ state dominates.
For the non-Peierls ECE-C and EHE-H degeneracy lines, we can’t use Pirogov-Sinai, so the following theorem is proven using disagreement percolation and translation invariance of the unique Gibbs state.
Theorem 6.
Along the ECE-C degeneracy line, if $e_{F}\geq e_{H}$ then there is a unique Gibbs state, for which no configuration dominates at any temperature. The same is true along the HFH-F degeneracy line when $e_{E}\geq e_{C}$.
We also obtain a weaker analogue of domination as an application of reflection positivity, which all Hadwiger models satisfy.
Theorem 7.
Along the ECE-C degeneracy line, if $e_{F}\geq e_{H}$ then probability a vertex is in either the $H$ or $F$ states decays exponentially in the inverse temperature. The same is true for $E$ and $C$ states along the HFH-F degeneracy line where $e_{E}\geq e_{C}$.
@@
Figure 5: Phase diagram showing the regions where given configurations dominate. Dark regions surrounding the non-Peierls degeneracy lines indicate areas where Pirogov–Sinai techniques are not applicable. This region shrinks with decreasing temperature, but is never empty. Outside of those areas each color, including white, represents a region where a particular set of configurations dominates. On the boundaries between these regions multiple Gibbs states not related by a symmetry dominate. On some of the non-Peierls lines we use disagreement percolation to prove no domination at any temperature.
3 Proofs of the the Main Theorems